Jump to main content.


Drake Chemical

EPA Daily Situation Reports for September 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, September 30, 1998

FOR DATES: From 7:00 a.m. Tuesday, Sept. 29, 1998 to 7:00 a.m. Wednesday, Sept. 30, 1998.

WEATHER: Partly cloudy. The temperature range was 44oF to 72oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 847.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.9 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the feed screws. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed at the southern end of the exclusion zone. The contractor performed a survey of the excavation bottom. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-09-25-98-232) for bin 2. Test results indicated that the thermally treated soil met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The TDF is currently scheduled to be off line sometime next week (tentatively October 6th and 7th ) to perform several maintenance tasks, including the replacement of the kiln feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Mike Welch & Sandy Downs (PADEP), Matt Radek (OHM)

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Tuesday, September 29, 1998

FOR DATES: From 7:00 a.m. Monday, Sept. 28, 1998 to 7:00 a.m. Tuesday, Sept. 29, 1998.

WEATHER: Partly cloudy and breezy. The temperature range was 44oF to 71oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 815.2 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the performance of the weekly weighbelt calibration and to measure the wear on the dual feed screws. The TDF processed contaminated soil at an average rate of 34.5 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the feed screws. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-09-23-98-230 and bot-ash-09-24-98-231) for bins 4 and 3 respectively. Test results indicated that the thermally treated soil met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The TDF is currently scheduled to be off line sometime next week (tentatively October 6th and 7th ) to perform several maintenance tasks, including the replacement of the kiln feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Monday, September 28, 1998

FOR DATES: From 7:00 a.m. Sunday, Sept. 27, 1998 to 7:00 a.m. Monday, Sept. 28, 1998.

WEATHER: Sunny and breezy w/ thunderstorms in the evening. The temperature range was 60oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 20.32 hours and 687.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the feed screws. The majority of offline time was attributed to the stoppage of feed during local thunderstorm activity. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operation, feed preparation, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling, and ash backfilling. A large volume of chain link fence encountered during excavation was decontaminated and placed into a rolloff for shipment to a permitted offsite disposal facility. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-09-22-98-229) and fly ash analytical report (flyash-09-22-98-043). Test results indicated that the bottom ash met project requirements for backfill on-site above elevation 547. Fly ash met project requirements for restricted backfill on-site above elevation 553 due to the concentrations of cadmium, chromium, and selenium being greater than the PADEP drinking water standard (DWS), but less than 25 times the DWS.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The TDF is currently scheduled to be off line sometime next week (tentatively October 6th and 7th ) to perform several maintenance tasks, including the replacement of the kiln feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for Dave Modricker (USACE)


Sunday, September 27, 1998

FOR DATES: From 7:00 a.m. Saturday, Sept. 26, 1998 to 7:00 a.m. Sunday, Sept. 27, 1998.

WEATHER: Party sunny. The temperature range was 59oF to 81oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.95 hours and 817.9 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time was due to the daily weigh belt calibration. The TDF processed contaminated soil at an average rate of 34.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, September 26, 1998

FOR DATES: From 7:00 a.m. Friday, Sept. 25, 1998 to 7:00 a.m. Saturday, Sept. 26, 1998.

WEATHER: Overcast with light rain. The temperature range was 55oF to 63oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.2 hours and 801.8 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of feed during local lightning storm activity. The TDF processed contaminated soil at an average rate of 34.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor also performed a site survey and poured the concrete base for the proposed lime silo. Excavation activities continued along the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-09-21-98-228) for bin 6. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, September 25, 1998

FOR DATES: From 7:00 a.m. Thursday, Sept. 24, 1998 to 7:00 a.m. Friday, Sept. 25, 1998.

WEATHER: Clear and cool. The temperature range was 38oF to 69oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 851.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.8 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the feed screws. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, debris handling, ash handling, and ash backfilling. Excavation activities were performed along the corridor sheetpiling on this date. MRI finished the preventative maintenance and leak checks on the heat trace lines of the perimeter air monitoring (PAM) system. The PAM stations were brought back on-line in the early afternoon. While the system was down, OHM's health and safety department performed continuous air monitoring with hand held devices.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Dave Mummert (OHM-CIH)

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Thursday, September 24, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 23, 1998 to 7:00 a.m. Thursday, Sept. 24, 1998.

WEATHER: Partly sunny and cooler. The temperature range was 38oF to 64oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 817.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.3 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the feed screws. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, debris handling, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor performed a survey of the 2-foot rock bottom. Excavation in sediment basin number 1 continued on this date. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Wednesday, September 23, 1998

FOR DATES: From 7:00 a.m. Tuesday, Sept. 22, 1998 to 7:00 a.m. Wednesday, Sept. 23, 1998.

WEATHER: Partly Cloudy and warm, light evening rain. The temperature range was 58oFto74oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 798.4 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.5 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the screws. Down time was due to weigh belt calibration, a feed screw trip, and a high kiln breech pressure. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, debris handling, ash handling, and ash backfilling. The contractor continued excavation on the southern face of the contaminated soil stockpile. Continued with the real time perimeter air monitoring (PAM). Received bottom ash analytical report (bot-ash-09-18-98-225) from bin 2. Bottom ash met project requirements for unrestricted backfilling on-site (i.e. above elevation 547). The contractor is forming a concrete footer pad for the installation of a lime silo.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Rick Lupold and David Modricker (USACE)


Tuesday, September 22, 1998

FOR DATES: From 7:00 a.m. Monday, Sept. 21, 1998 to 7:00 a.m. Tuesday, Sept. 22, 1998.

WEATHER: Partly Cloudy and warm, light evening rain. The temperature range was 65oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 843.4 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.6 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the screws. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, debris handling, ash handling, and ash backfilling. The contractor continued excavation on the southern face of the contaminated soil stockpile. Continued with the real time perimeter air monitoring (PAM). Received bottom ash analytical report (bot-ash-09-17-98-224) from bin 3. Bottom ash met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE)


Monday, September 21, 1998

FOR DATES: From 7:00 a.m. Sunday, Sept. 20, 1998 to 7:00 a.m. Monday, Sept. 21, 1998.

WEATHER: Warm, Hazy, Rain overnight. The temperature range was 61oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.77 hours and 869.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 36.6 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the screws. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation and handling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The contractor is evaluating the wear on the feed screws since the addition of cutting teeth on Thursday, September 17, 1998. The contractor has no down time scheduled in the near future to replace the feed screws and perform TDF system wide maintenance.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: David Modricker (USACE)


Sunday, September 20, 1998

FOR DATES: From 7:00 a.m. Saturday, Sept. 19, 1998 to 7:00 a.m. Sunday, Sept. 20, 1998.

WEATHER: Warm, Hazy, Humid. The temperature range was 59oF to 75oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.95 hours and 849.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris separation and handling, ash handling, and ash backfilling. The contractor applied Concover to the previously excavated sediment basin 1 area. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, September 19, 1998

FOR DATES: From 7:00 a.m. Friday, Sept. 18, 1998 to 7:00 a.m. Saturday, Sept. 19, 1998.

WEATHER: Clear and mild. The temperature range was 51oF to 81oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 841.7 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Site soil and lime were mixed with the soft sediments of basin 1. The soil was later excavated from basin 1 for thermal treatment. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, September 18, 1998

FOR DATES: From 7:00 a.m. Thursday, Sept. 17, 1998 to 7:00 a.m. Friday, Sept. 18, 1998.

WEATHER: Light rain in the morning, clearing later. The temperature range was 51oF to 75oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 14.4 hours and 515.3 tons of contaminated soil were thermally treated during the time frame of this report. Feed was shut down at 0745 hours in order to perform the monthly TDF interlock checks and calibrations and to perform maintenance on the feed screws. The TDF processed contaminated soil at an average rate of 35.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor performed TDF operations, interlock checks, and maintenance to the dual feed screws. As part of the interlock checks, the Thermal Relief Valve (TRV) was opened at 1030 hours. Feed to the incinerator had been shut down for approximately 3 hours prior to the opening. Before the TRV was opened, the contractor notified the proper regulatory authorities. The contractor also continued WWTP operations, feed preparation, lime mixing, wood shredding, debris handling, debris separation, ash handling, and ash backfilling. Excavation continued on the western end of sediment basin 1 and dry soil and lime was mixed with the soft sediment of basin 1. Continued with the real time perimeter air monitoring (PAM).

Received bottom ash analytical report (bot-ash-09-13-98-220) for bin 1. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-09-13-98-041) for bin D. Test results indicated that the fly ash materials met project requirements for restricted backfilling (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Planned maintenance on the TDF is scheduled in the upcoming weeks.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for Frank Cleaver (USACE)


Thursday, September 17, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 16, 1998 to 7:00 a.m. Thursday, Sept. 17, 1998.

WEATHER: Hazy with heavy rain in the afternoon. The temperature range was 65oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 951.5 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

Feed to the TDF will be stopped Thursday morning, September 17, 1998 to perform monthly interlock checks and instrument calibrations. There will be a planned Thermal Relief Valve (TRV) opening as part of the interlock check, and appropriate notifications will be made. During the period that feed is shut off, the contractor will modify the soil feed screws to the TDF, including installing cutting teeth on the screws. The planned shut down is scheduled to last approximately 8 hours.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, dust control, debris handling, debris separation, ash handling, and ash backfilling. Excavation was continued on the southwest corner of the excavation face. Continued with the real time perimeter air monitoring (PAM) program. AirRecon, the stack testing subcontractor, performed the final run of the quarterly stack testing for metals and particulate. Received bottom ash analytical report (bot-ash-09-12-98-219) for bin 7. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Planned maintenance on the TDF is scheduled in the upcoming weeks.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Chris McBride (FOCUS Environmental), Dr. Scott Tucker (OHM), Kurt Davis (AIR)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, September 16, 1998

FOR DATES: From 7:00 a.m. Tuesday, Sept. 15, 1998 to 7:00 a.m. Wednesday, Sept. 16, 1998.

WEATHER: Hot and Hazy. The temperature range was 64oF to 91oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for approx. 24 hours and 945.7 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation. AirRecon is on-site to perform the quarterly metals and particulate stack testing.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, dust control, debris handling, debris separation, ash handling, and ash backfilling. The contractor performed a survey of the excavation and began excavation on the southwest corner of the excavation face. Continued with the real time perimeter air monitoring (PAM) program. AirRecon, the stack testing subcontractor, performed the first and second run of the quarterly stack testing for metals and particulate.

Received bottom ash analytical report (bot-ash-09-11-98-218) for bin 2. Test results indicated that all of the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Carmine Leone & Benny Hom (USACE), Chris McBride (FOCUS Environmental), Dr. Scott Tucker (OHM), Sandy Downs & Steve Webster (PADEP), Kurt Davis (AIR)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, September 15, 1998

FOR DATES: From 7:00 a.m. Monday, Sept. 14, 1998 to 7:00 a.m. Tuesday, Sept. 15, 1998.

WEATHER: Hot and Hazy. The temperature range was 63oF to 89oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 24 hours and 964.7 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation. AirRecon is on-site to perform stack testing for metals and particulate.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with the real time perimeter air monitoring (PAM) program and finished the special sampling for metals and particulate at the PAM stations. AirRecon, the stack testing subcontractor, mobilized to the site.

Received bottom ash analytical reports (bot-ash-09-10-98-217, bot-ash-09-08-98-215, bot-ash-9-9-98-216) for bin 5, bin 3, and bin 4 respectively. Test results indicated that all of the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Chris McBride (FOCUS Environmental), Kurt Davis (AIR)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, September 14, 1998

FOR DATES: From 7:00 a.m. Sunday, Sept. 13, 1998 to 7:00 a.m. Monday, Sept. 14, 1998.

WEATHER: Warm and Hazy. The temperature range was 62oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.5 hours and 908.3 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time is attributed to the replacement of a drive chain link for the rotary kiln. The TDF processed contaminated soil at an average rate of 40.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, and ash handling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-09-05-98-212) for bin 6. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The stack testing subcontractor is expected to arrive on site Monday, September 14, 1998. Actual stack testing for metals and particulates is scheduled to start on September 15, 1998 and continue on September 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Sunday, September 13, 1998

FOR DATES: From 7:00 a.m. Saturday, Sept. 12, 1998 to 7:00 a.m. Sunday, Sept. 13, 1998.

WEATHER: Mostly clear with increasing humidy. The temperature range was 55oF to 89oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.85 hours and 966.5 tons of contaminated soil were thermally treated during the time frame of this report. Down time was attributed primarily to conducting a weigh belt calibration. The TDF processed contaminated soil at an average rate of 40.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Greg Beauseigneur for David Modricker (USACE)


Saturday, September 12, 1998

FOR DATES: From 7:00 a.m. Friday, Sept. 11, 1998 to 7:00 a.m. Saturday, Sept. 12, 1998.

WEATHER: Partly cloudy, clearing in the afternoon. The temperature range was 47oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 18.0 hours and 703.2 tons of contaminated soil were thermally treated during the time frame of this report. Down time was attributed to the replacement of a broken coupling that connected the feed screw motor to the feed screws. Additional down time was attributed to the stoppage of feed to facilitate the transportation of an injured worker from the exclusion zone. The TDF processed contaminated soil at an average rate of 39.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program and began the OHSA method 93 24-hour air sampling event for BNA at the off-site air monitoring stations and at the PAM stations.

Received bottom ash analytical reports (bot-ash-09-06-98-213 and bot-ash-09-07-98-214) for bins 7 and 1 respectively. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical reports (fly-ash-09-04-98-039 and fly-ash-09-07-98-040) for bins B and C respectively. Test results indicated that both fly ash samples met project requirements for restricted backfilling on-site (i.e. greater than elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: An injured worker was transported from the site to the Lock Haven Hospital due to contusions on the buttocks and back. The worker misjudged the distance to a landing, and fell while descending steps in the feed prepartion building which is in the exclusion zone. The worker was later released from the hospital and returned to work the same shift.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, September 11, 1998

FOR DATES: From 7:00 a.m. Thursday, Sept. 10, 1998 to 7:00 a.m. Friday, Sept. 11, 1998.

WEATHER: Clear and breezy. The temperature range was 46oF to 74oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 966.3 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, stockpile grading, lime mixing, wood shredding, debris handling, debris separation, debris washing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Concover, an odor and erosion suppressant, was placed on the outer slopes of the ash stockpile. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson and David Modricker (USACE)


Thursday, September 10, 1998

FOR DATES: From 7:00 a.m. Wednesday, Sept. 9, 1998 to 7:00 a.m. Thursday, Sept. 10, 1998. WEATHER: Partly cloudy and windy. The temperature range was 44oF to 66oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 966.2 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, stockpile grading, lime mixing, wood shredding, debris handling, debris separation, debris washing, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE)


Wednesday, September 9, 1998

FOR DATES: From 7:00 a.m. Tuesday, Sept. 8, 1998 to 7:00 a.m. Wednesday, Sept. 9, 1998.

WEATHER: Partly cloudy with periods of rain. The temperature range was 44oF to 67oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.0 hours and 877.3 tons of contaminated soil were thermally treated during the time frame of this report. Down time is attributed to the manual stoppage of feed during local lightning storm activity, and to an automatic feed stoppage as a result of the wet ash drag conveyor backing up with ash materials at the ash receiving bin. The TDF processed contaminated soil at an average rate of 39.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, debris washing, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-09-03-98-210 and bot-ash-09-02-98-209) for bins 5 and 4 respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: As a precautionary measure, feed to the TDF automatically shut down at 0320 hours due to an excess of bottom ash materials collecting in the ash receiving bin and backing up to the wet ash drag conveyor. The ash receiving bin was immediately emptied, the wet ash drag was cleared of excess material, and feed was restarted approximately 1 hour later.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson and David Modricker (USACE)


Tuesday, September 8, 1998

FOR DATES: From 7:00 a.m. Monday, September 7, 1998 to 7:00 a.m. Tuesday, September 8, 1998.

WEATHER: Clear and mild. The temperature range was 50oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.15 hours and 894.3 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.4 tons per hour while on line. Feed was stopped late in the morning as a precaution due to local thunderstorm activity. During the down time, the daily weigh belt calibration was conducted and the feed screws were measured to estimate the amount of wear. Contaminated soil feed resumed shortly after noon. Down time was also attributed to a flame out in the Secondary Combustion Chamber burner. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-09-02-98-209). Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: David Modricker (USACE)


Monday, September 7, 1998

FOR DATES: From 7:00 a.m. Sunday, Sept. 6, 1998 to 7:00 a.m. Monday, Sept. 7, 1998.

WEATHER: Clear and mild. The temperature range was 60oF to 90oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.88 hours and 969.9 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.6 tons per hour while on line. Down time was necessary to conduct a daily weigh belt calibration. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: David Modricker (USACE)


Sunday, September 6, 1998

FOR DATES: From 7:00 a.m. Saturday, Sept. 5, 1998 to 7:00 a.m. Sunday, Sept. 6, 1998.

WEATHER: Clear and mild. The temperature range was 50oF to 81oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.65 hours and 958.7 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.5 tons per hour while on line. Down time was due to a weigh belt calibration and a flame out in a kiln burner. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, September 5, 1998

FOR DATES: From 7:00 a.m. Friday, Sept. 4, 1998 to 7:00 a.m. Saturday, Sept. 5, 1998.

WEATHER: Clear and cool. The temperature range was 51oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 967.5 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning and backfilling, wood shredding, ash handling, and ash backfilling. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-08-29-98-205, bot-ash-08-31-98-207, and bot-ash-08-30-98-206) for bins 6, 1, and 7, respectively. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-08-30-98-038) for bin A. Test results indicate that the fly ash materials met project specifications for restricted backfilling (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Dick Lee (USACE)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, September 4, 1998

FOR DATES: From 7:00 a.m. Thursday, Sept. 3, 1998 to 7:00 a.m. Friday, Sept. 4, 1998.

WEATHER: Overcast and clearing later. The temperature range was 53oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 968 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, stockpile grading, feed preparation, lime mixing, debris handling, debris separation, rock cleaning and backfilling, wood chipping, ash handling, and ash backfilling. Concover, an odor suppressant and erosion controlling material, was placed on the recently graded stockpile. The contractor performed no new excavation on the southern face of the stockpile on this date. Continued with perimeter air monitoring (PAM) program, and began the 24-hour NIOSH 5518 and OSHA method 93 testing for BNA at the off-site and on-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, September 3, 1998

FOR DATES: From 7:00 a.m. Wednesday, September 2, 1998 to 7:00 a.m. Thursday, September 3, 1998.

WEATHER: Overcast with afternoon rain. The temperature range was 54oF to 73oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 965.6 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, stockpile grading, feed preparation, lime mixing, debris handling, debris separation, rock cleaning and backfilling, wood chipping, ash handling, and ash backfilling. The contractor performed no excavation on the southern face of the stockpile on this date. Received bottom ash analytical report (bot-ash-08-25-98-201) for bin 2. Test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Continued with perimeter air monitoring (PAM) program. Began utilizing the backup real time air monitoring system (HNu 201 and SRI gas chromatograph) while the primary real time air monitoring system (HNu 501) is being repaired.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, September 2, 1998

FOR DATES: From 7:00 a.m. Tuesday, September 1, 1998 to 7:00 a.m. Wednesday, September 2, 1998.

WEATHER: Sunny and clear. The temperature range was 52oF to 79oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 963.6 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, stockpiling, feed preparation, lime mixing, debris handling, debris separation, oversized rock cleaning and backfilling, wood chipping, ash handling, and ash backfilling. The contractor performed a monthly ash volume survey of the ash stockpile. Continued real time perimeter air monitoring and began the monthly time weighted average perimeter air sampling event.

Received bottom ash analytical reports (bot-ash-08-27-98-203 and bot-ash-08-28-98-204) for bins 4 and 5 respectively. Test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: S. Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, September 1, 1998

FOR DATES: From 7:00 a.m. Monday, Aug. 31, 1998 to 7:00 a.m. Tuesday, September 1, 1998.

WEATHER: Sunny and clear. The temperature range was 52oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 955.3 tons of contaminated soil was thermally treated during the time frame of this report. Down time is attributed to a conducting a weigh belt calibration and to check the wear on the new feed screws. The TDF processed contaminated soil at an average rate of 40.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, stockpiling, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, wood chipping, ash handling, and ash backfilling. Continued real time perimeter air monitoring. The contractor held a site orientation for new hires that had recently completed the required OSHA 40-hour training for hazardous waste workers.

Received fly ash analytical report (fly-ash-08-26-98-037) for bin E. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on site (i.e. above elevation 553) due to the concentrations of lead, selenium, and chromium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulates is scheduled for September 14 through 16, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.